Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (ending 12 Sep); Correlator hardware (62 BlB pairs) available. Double maintenance day with power cyclings/fire suppression testing (=disruptive); unlikely sci-ops until Thursday.
  • Key dates:
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
  • Staff Travel:
    • FO: 12 Aug - 12 Sep
    • RP: 10 Sep - 16 Sep (Brazil)
    • URSI (Perley)
  • Visitors:
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)
    • Dave Roberts: 10 Aug - 20 Aug (SS433 Demo science)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA25 out; EA23 on MP; back to array 1 Sep. Problems over the weekend with EA27 (subreflector); EA11 power;
    • Receivers: EA23; EA24 for first L-band solar capable Rx.
      • Perley: RF Switch testing ongoing. Meeting Wednesday; AOC 317 3pm MDT (Jackson/Scott).
    • Calibration (Perley): Gain compression/expansion ongoing (Hayward).
    • Brentjens: More testing this week.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): Some issues with BlB prs over the weekend suspiciously related to power glitches. Work on 3-bit sampler auto-gain setting this week.
      • Review of last week's items: 13-17 Aug: 1E37 amps (CBE bad initialization of no data); 15-16 Aug: high frequency problems (bad pointing script); problem interframe delay (improved transition testing); malformed subset of BDFs (no clear; test in place - no recurrence).
  • Computing (Robnett):
    • Notes: Issues with lustre filling - please move/remove data as possible; a new lustre node is in and likely to be installed this week.

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon
    • 1000-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Identified possible issues in the noise related to the 3-level phase rotator. Water maser data taken. Write-up underway to be shared with Brent; further planning pending
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). .Needs close-out testing (Michael); confirmation of visibility/flags (Joe).
  • Correlator: [Clark/Carlson] Delay rocking across bands. Ongoing discussion of implementation; pending Bryan's return
    • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).
    • Array: Hunter: OPT testing revealed some oddities with North Arm antennas (status)?
    • Array: Chandler: Stepped (asymptotic) gain changes on some antenna IFs (status)?
    • Array: Hughes: Rapid phase changes at L band
    • Array: EA28 many zeros
    • Array: EA24 A/C (- bad L301-1 (wouldn't lock)-being replaced today).
    • Array: EA27 L305 instability; being looked at today.

Testing This week/soon

  • [Sowinski et al]; Executor update testing
    • Phased array testing as possible
  • Planetary tracking tests
  • Holography scanning tests
  • 3-bit sampler testing; automatic gain setting.
  • astrometry mode testing.
  • MCAF update testing.
  • [Goss et al]; Sub-band edge effects testing.
  • [Owen]: L band deep field study
  • Subarrays: cleanup.
  • Stage 2 mixer testing to relax sub 128 MHz BW grids
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Roy/Dhawan: HI High resolution bandpass testing
  • Baseband edge/filter rolloff.
  • Sub-band dependent fshifts.
  • Array: *; Individual issue sleuthing.

Commissioning Groups

    • Adv: Low band (Owen):
      • Notes:
    • Adv: Astrometry (Mioduszewski):
      • Notes: Initial UVW tests: I wanted to confirm that CASA and AIPS are close (~few parts in 10^6). In fact, they are not. UVFIX run on data imported from CASA makes a change of almost 10^-4.
  • Calibrator Models
    • Notes: Reduction underway.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes: ~500 GB data set obtained.
  • Imaging (Ott/Golap):
    • Notes: Takahashi/Ott working on mosaic needs in PST/OPT. Document under review.
  • Algorithm (Bhatnagar, Owen):

RSRO/ECSO Reports

RSRO A configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

  • 11A-142 (Hughes): EA28 issues identified; phase drifts over time during day time observing (L).

Misc:

Demo Science Status

AOB

-- JosephMcMullin - 2011-08-23
Topic revision: r2 - 2011-08-29, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback